OpenStudy (anonymous):

convert 25 yards to inches

OpenStudy (triciaal):

1 yard = 36 inches 25 * 1 yd = 25 * 36 inches

OpenStudy (anonymous):

Theres 3 feet in a yard, and 12 inches in a foot. So 3*25=x than x*12 is your answer.